OpenStudy (anonymous):

what is the equation of a line tangent to y=x^2?

OpenStudy (anonymous):

depends on the point the equation will be different for different points on the curve

OpenStudy (anonymous):

for example, at \((3,9)\) it will be \[y-9=6(x-3)\]

OpenStudy (anonymous):

oh i the problem was asking for what k equals in the equation y=kx to have 1 solution the the equation y=x^2. i interpreted it, probably wrongly, as it asking for a tangent line

OpenStudy (anonymous):

i don't understand the problem what is \(k\) so that the equation of \(y=kx\) will be TANGENT

OpenStudy (anonymous):

we are given the graph of y=x^2. we have to find an equation of a line in the form y=x+k so that there i 1 solution to this system of equations. what value is k when there is 1 solution the the system?

OpenStudy (anonymous):

It is a tangent line. And if you KNOW it is y=x+k then you have the slope. 1. So find the derivative: D(x^2)=2x=1 implies x=1/2. So now you have (1/2,1/4) as your point and 1/2 as your slope. y-y_0=m(x-x_0) you're done.